### v1

# Reinforcement Fine-Tuning Naturally Mitigates Forgetting in Continual Post-Training

Abstract
Continual post-training (CPT) is a popular and effective technique for adapting foundation models like multimodal large language models to specific and ever-evolving downstream tasks. While existing research has primarily concentrated on methods like data replay, model expansion, or parameter regularization, the fundamental role of the learning paradigm within CPT remains largely unexplored. This paper presents a comparative analysis of two core post-training paradigms: supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT), investigating their respective impacts on knowledge retention during CPT. Our experiments are conducted on a benchmark comprising seven diverse multimodal tasks, utilizing Qwen2.5-VL-7B-Instruct as the base model for continual post-training. The investigation yields two significant findings: (1) When continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks. In contrast, RFT inherently preserves prior knowledge and achieve performance comparable to multi-task training. (2) RFT successfully protects and even enhances the model’s general knowledge on standard benchmarks (e.g., MMMU and MMLU-Pro). Conversely, SFT degrades general model capabilities severely. Further analysis shows that explicit mechanisms, such as KL penalty and chain-of-thought reasoning, are not the primary factors. Instead, we find that the implicit regularization inherent to RFT is a key factor in mitigating forgetting. Finally, we propose a rollout-based instance filtering algorithm to improve the stability and efficiency of RFT. Our comprehensive study demonstrates the superiority of RFT as a robust paradigm for continual post-training.
Figure 1: Comparison of performance retention between SFT and RFT in continual post-training. We plot the performance relative to the initial training performance on each task as learning progresses through a sequence of tasks. (a) SFT exhibits catastrophic forgetting; (b) RFT maintains performance on previously learned tasks.
1 Introduction
Recent advancements in multimodal large language models (MLLMs) have demonstrated remarkable capabilities in complex world understanding (Achiam et al., 2023; Liu et al., 2024; Wang et al., 2024a). To align with the demands of real-world deployment, MLLMs must adapt to a stream of data and evolving user requirements, incorporating new skills and domain knowledge over time (Zhu et al., 2024). This calls for an efficient and scalable continual post-training (CPT) paradigm. A key challenge in CPT is the well-known phenomenon of catastrophic forgetting (McCloskey & Cohen, 1989), where adapting to a new task leads to a severe degradation of performance on previously learned tasks. To reduce forgetting, recent studies (Guo et al., 2025c) focus on data replay (Maharana et al., 2025; Lee et al., 2025; Wang et al., 2025), model expansion (Zhao et al., 2025; Guo et al., 2025b; Zeng et al., 2024), and explicit regularization (Liu et al., 2025a). Nevertheless, existing methods typically leverage the supervised fine-tuning (SFT) paradigm by default, and the role of the fundamental fine-tuning paradigm in CPT has been overlooked.
Recently, reinforcement fine-tuning (RFT), which optimizes models based on feedback from generated outputs, has significantly advanced foundation model post-training (Chu et al., 2025; Shao et al., 2024; Guo et al., 2025a). In this work, we conduct the first investigation into whether SFT or RFT is the more suitable paradigm for CPT, focusing on knowledge preservation for both specific downstream tasks and general capabilities. Experimentally, we continually fine-tune the Qwen2.5-VL-7B-Instruct model (Bai et al., 2025b) on a benchmark comprising diverse multimodal tasks covering various domains. To fully reflect the knowledge preservation ability, we evaluate forgetting on both learned specific tasks and general benchmarks such as MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a).
The empirical investigation yields two notable findings: (1) As shown in Figure 1, when continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks, which is consistent with existing studies (Guo et al., 2025c). In contrast, RFT can inherently protect prior knowledge, maintaining strong performance on old tasks after being adapted to new tasks. Surprisingly, without any data replay, continual post-training with RFT can achieve comparable performance with that of multi-task training, which is not achievable even when equipping SFT with continual learning strategies. (2) As demonstrated in Figure 2, continual training on downstream tasks with SFT severely degrades general model capabilities, which is known as base model degradation (Liu et al., 2025a). For example, the performance drops from 52.1% to 40.1% on MMMU. Fortunately, RFT protects the general performance and enhances the model’s general knowledge (52.1% → 54.2%). These observations highlight the knowledge preservation capability of RFT.
Figure 2: General capability preservation after continual post-training. We evaluate models at the end of learning all downstream tasks on general benchmarks using both CoT and direct prompting. Compared to the base model, SFT (shown in light colors) causes degradation while RFT (shown in darker colors) preserves and even enhances general capabilities.
To understand how RFT mitigates forgetting during CPT, we conduct additional experiments with the popular and representative group relative policy optimization (GRPO) framework (Shao et al., 2024). We analyze the impact of KL divergence penalty and chain-of-thought (CoT) reasoning (Wei et al., 2022) on forgetting mitigation. Particularly, the KL divergence penalty prevents the policy from changing too drastically, similar to the well-known knowledge distillation in continual learning (Li & Hoiem, 2017). However, we reveal that KL penalty and CoT mechanisms are not the primary factors, and instead, the implicit regularization introduced by RFT significantly contributes to forgetting mitigation. Concretely, the expected magnitude of the gradient update, when measured in the parameter subspace sensitive to old tasks, is inherently more conservative than that of SFT. This conservatism is naturally scaled by the variance of the reward signal, effectively creating a data-dependent regularization. Last but not least, we observe that the learning process of RFT can be highly inefficient. Thus, we introduce a rollout-based instance filtering algorithm that enhances the stability of GRPO while still being an excellent knowledge protector.
Our main contributions are threefold:
  1. 1.
    We present a comprehensive analysis of the forgetting mitigation effects of SFT and RFT during continual post-training of MLLMs, demonstrating that RFT naturally preserves not only the performance of learned downstream tasks but also general model capabilities.
  2. 2.
    Based on in-depth analyses, we reveal that the implicit regularization introduced by RFT significantly contributes to the forgetting mitigation, being more important than KL regularization and CoT reasoning.
  3. 3.
    We propose a rollout-based instance filtering algorithm that enhances the stability and efficiency of RFT while still maintaining previous learned knowledge.
2 Related Works
Continual Post-Training in MLLMs.
Continual learning aims to enable models to learn from a stream of tasks without catastrophically forgetting previously acquired knowledge (Van de Ven et al., 2022). For MLLMs, this capability is particularly important for adapting these powerful models to a diverse range of downstream multimodal tasks. Existing CPT research in MLLMs (Guo et al., 2025c) has focused on adapting traditional forgetting mitigation strategies such as regularization, data replay, and model expansion, within an SFT paradigm. Regarding benchmark, Chen et al. (2024) introduced a continual instruction tuning benchmark including several specific multimodal datasets. Zhao et al. (2025) introduces two settings named domain continual learning and ability continual learning, providing a realistic evaluation for continual post-training of MLLMs. In addition to these methods, recent efforts to mitigate catastrophic forgetting in MLLMs primarily focus on parameter-efficient learning and dynamic data selection. For instance, HiDe-LLaVA (Guo et al., 2025b) employs a hierarchical decoupling framework for task-specific LoRA expansion and general knowledge fusion. MRLoRA (Zhao et al., 2025) leverages architectural decoupling and a multimodal routing mechanism to selectively activate specialized parameters. In terms of data management, Adapt-$\infty$ (Maharana et al., 2025) dynamically selects high-impact samples based on gradient representations and prunes redundant data. These diverse strategies collectively aim to enhance the ability of MLLMs to continually learn new tasks while preserving previously acquired knowledge. Recently, Liu et al. (2025a) developed LLaVA-c, which is a simple yet effective CPT framework for MLLMs, addressing task balancing and catastrophic forgetting through spectral-aware consolidation and unsupervised inquiry regularization.
Post-Training of Foundation Models.
Post-training is a critical stage for refining the capabilities of pre-trained foundation models (Shao et al., 2024; Chu et al., 2025; Achiam et al., 2023). SFT on task-specific or instruction-formatted datasets is a common approach to adapt models to downstream applications (Chung et al., 2024; Zhou et al., 2023). For example, Chung et al. (2024) demonstrated that by scaling the number of tasks and model size, and incorporating CoT data, SFT significantly enhances the performance and generalization of various large language models across diverse benchmarks. Recently, RFT has gained prominence for aligning models with human preferences or improving performance on specific objectives (Liu et al., 2025c; Zhai et al., 2024; Shao et al., 2024; Luong et al., 2024; Li et al., 2025; 2023c; Ahmadian et al., 2024). Particularly, GRPO (Shao et al., 2024) largely enhances mathematical reasoning and optimizes memory usage, being a popular method for post-training of large language models. Liu et al. (2025b) revealed inherent biases in the GRPO algorithm, then introduces an unbiased optimization method that improves token efficiency while maintaining reasoning performance. Visual-RFT (Liu et al., 2025c) boosts MLLMs by using reinforcement learning with rule-based visual rewards, making them more data-efficient and better at various visual tasks than traditional SFT. Recently, Chu et al. (2025) demonstrated that reinforcement learning significantly enhances the generalization capabilities of foundation models, while SFT primarily leads to memorization. In this work, we study the comparative effect of SFT and RFT on knowledge retention in MLLMs continual post-training.
3 Preliminaries
Post-training is a critical phase following large-scale pre-training that adapts foundation models to specific downstream tasks or align them with human preferences (Ouyang et al., 2022; Kumar et al., 2025). We model the MLLM with parameters $\theta$ as a policy $\pi_{\theta}$. This policy defines a conditional probability distribution $\pi_{\theta}(a|x)$ over possible text responses $a$ given a multimodal input prompt $x$, which consists of text and images. We also assume a scalar reward function $r(x,a)\in\mathbb{R}$ that evaluates the quality of a response. Post-training aims to update the parameters $\theta$ of a pre-trained base model $\pi_{\theta_{\text{base}}}$ to improve its performance on a downstream task using a training dataset $\mathcal{D}$, which can be achieved by SFT (Ouyang et al., 2022) or RFT (Lee et al., 2023).
SFT.
Given training dataset $\mathcal{D}=\{(x_{i},a_{i}^{*})\}_{i=1}^{N}$ consisting of prompts $x_{i}$ and their corresponding ground-truth responses $a_{i}^{*}$, SFT maximizes the likelihood of generating the ground-truth responses. This is typically achieved by minimizing the negative log-likelihood loss:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(x,a^{*})\sim\mathcal{D}}[\log\pi_{\theta}(a^{*}|x)]=-\mathbb{E}_{(x,a^{*})\sim\mathcal{D}}\left[\sum_{t=1}^{|a^{*}|}\log\pi_{\theta}(a_{t}^{*}|x,a_{<t}^{*})\right].$ (1)
RFT.
In RFT, the model $\pi_{\theta}$ is treated as a policy, and generates one or more candidate responses for a given prompt $x$. The optimization objective is to maximize the expected reward:
$\mathcal{J}_{\text{RFT}}(\theta)=\mathbb{E}_{x\sim\mathcal{D}}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}[r(x,a)].$ (2)
The gradient of this objective is typically estimated using policy gradient methods. The most basic form is the REINFORCE (Williams, 1992) estimator, which, unfortunately, has high gradient variance. Recent RFT algorithms (Shao et al., 2024; Li et al., 2023c; Ahmadian et al., 2024) address this issue by designing more stable advantage estimators and baselines. We introduce some of the representative methods used in our study below.
For a prompt $x$, GRPO (Shao et al., 2024) generates a group of $n$ responses $\{a_{1},\dots,a_{n}\}$ and computes their rewards $\{r_{1},\dots,r_{n}\}$. The advantage for a response $a_{i}$ is its normalized reward relative to the group mean: $A(a_{i})=(r_{i}-\bar{r})/\sigma_{r}$, where $\bar{r}$ and $\sigma_{r}$ are the mean and standard deviation of the rewards. The objective is to maximize the expected advantage-weighted log-probability, often with a KL-divergence penalty against a reference policy $\pi_{\text{ref}}$ to stabilize training:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{a_{i}\}}\left[\sum_{i=1}^{n}A(a_{i})\log\pi_{\theta}(a_{i}|x)\right]-\beta D_{\text{KL}}(\pi_{\theta}(\cdot|x)||\pi_{\text{ref}}(\cdot|x)),$ (3)
where $\beta>0$. ReMax (Li et al., 2023c) use the reward of a greedy decoding response $a_{\text{greedy}}$ as a baseline. For a single sampled response $a$, the objective is to maximize:
$\mathcal{J}_{\text{ReMax}}(\theta)=\mathbb{E}_{x,a\sim\pi_{\theta}}\left[(r(x,a)-r(x,a_{\text{greedy}}))\log\pi_{\theta}(a|x)\right].$ (4)
This adaptive baseline helps to normalize rewards and reduce gradient variance. To further reduce variance, RLOO (Ahmadian et al., 2024) generates $n$ samples $\{a_{1},\dots,a_{n}\}$ and uses the average reward of the other $n-1$ samples as a baseline for sample $a_{i}$:
$\mathcal{J}_{\text{RLOO}}(\theta)=\mathbb{E}_{x,\{a_{i}\}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(r(x,a_{i})-\frac{1}{n-1}\sum_{j\neq i}r(x,a_{j})\right)\log\pi_{\theta}(a_{i}|x)\right].$ (5)
Continual Post-Training Formulation.
In CPT, the model learns from a sequence of $T$ tasks with datasets $\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$. The core challenge is catastrophic forgetting, i.e., a significant drop in performance on previously learned tasks. Following the general continual learning framework, CPT can be formulated as a constrained optimization problem. When learning task $t$, the objective is:
$\theta^{t}=\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{D}_{t})\quad\text{s.t.}\mathcal{L}(\theta;\mathcal{D}_{i})\leq\mathcal{L}(\theta^{i};\mathcal{D}_{i}),\quad\forall i\in[1,t-1]$ (6)
where $\mathcal{L}(\theta;\mathcal{D}_{i})$ is the training objective (e.g., negative log-likelihood for SFT or negative expected reward for RFT) on task $i$, and $\theta^{i}$ are parameters after learning task $i$.
4 Reinforcement Fine-Tuning Mitigates Forgetting in CPT
This section presents our comparative results comparing RFT and SFT in a continual post-training scenario. We detail our experimental setup and then present the main findings that highlight the superiority of RFT for knowledge preservation.
Figure 3: Illustrative examples of continual post-training benchmark.
4.1 Experimental Setup
Continual Post-Training Model & Datasets.
We adopt the open-source Qwen2.5-VL-7B-Instruct (Bai et al., 2025a) as our base model, primarily due to its demonstrated superiority in vision-language comprehension and its favorable resource footprint, which is crucial for practical deployment. We continually fine-tune the model on diverse vision-language datasets (ScienceQA (Saikh et al., 2022), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), GQA (Hudson & Manning, 2019), Geometry3K (Lu et al., 2021), PathVQA (He et al., 2020), Super-CLEVR (Li et al., 2023b)), covering a wide range of common downstream applications. After the end of CPT, evaluation is performed on the test sets of all previously encountered tasks. Additionally, to fully assess the knowledge preservation ability, we evaluate the model on diverse, general benchmarks at the end of learning all downstream tasks. Specifically, we evaluate the model on three specialized benchmarks: MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a). Particularly, we include POPE to systematically assess whether CPT induces object hallucination in MLLMs. A detailed description of those datasets is provided in the Appendix.
Learning Algorithms & Reward.
Our experiments encompass a range of fine-tuning algorithms, including standard SFT (Zheng et al., 2024) and several representative RFT algorithms, i.e., GPRO (Shao et al., 2024), ReMax (Li et al., 2023c), and RLOO (Ahmadian et al., 2024). For both SFT and RFT, model outputs are normalized by disregarding extraneous whitespace (e.g., spaces, indentations, newlines) and ignoring case sensitivity to ensure precise assessment. For GRPO, the overall reward $r_{\text{overall}}$ is designed with a weighted sum of accuracy reward and format reward:
$r_{\text{overall}}=0.9r_{\text{acc}}+0.1r_{\text{format}}.$ (7)
[…]
Prompt Template.
[…]
Evaluation Metrics.
To quantify the model’s performance during CPT, we adopt two standard metrics. Let $P_{t,j}$ denote the test accuracy on task $j$ after learning task $t$. We measure the final overall performance using average accuracy (AvgAcc), which is the average accuracy across all tasks after training on the final task $T$. To measure knowledge retention, we use the forgetting measure (FM), which calculates the average difference between the final accuracy of a task and the best accuracy achieved for that task throughout the training sequence. Let $P_{i}^{*}=\max_{k\in\{i,\dots,T\}}P_{k,i}$ be the best performance for task $i$. The above two metrics are defined as:
$AvgAcc=\frac{1}{T}\sum_{i=1}^{T}P_{T,i},\qquad FM=\frac{1}{T}\sum_{i=1}^{T}(P_{T,i}-P_{i}^{*}).$ (8)
A higher AvgAcc indicates better overall performance, while an FM closer to zero signifies less forgetting and better knowledge preservation.
Implementation Details.
All experiments employ full-parameter fine-tuning for both SFT and RFT to ensure comprehensive capability assessment. Experiments of SFT are conducted using the llamafactory (Zheng et al., 2024) framework, with a learning rate of $51e-$ 5 and a batch size of 24. RFT methods (GRPO, ReMax, and RLOO) are implemented using the easyR1 (Zheng et al., 2025) framework, building upon Verl (Sheng et al., 2024). A consistent configuration is applied across RFT methods to ensure an equitable comparison: a learning rate of $61e-$ 6, a rollout batch size of 512, a sampling temperature of 1.0, with KL-divergence coefficient $\beta=0.01$. Specifically, GRPO is implemented adhering to its foundational methodology, with a group size set to 8. ReMax followed its core algorithm, and RLOO adopted the official Hugging Face algorithm.
Table 1: Final performance comparison on all tasks after the entire continual learning sequence. The best and second-best results are highlighted. “-” indicates that the metric is not applicable.
Method SciQA TextVQA VizWiz GQA Geo. PathVQA sCLEVR AvgAcc FM Base 90.5 62.8 45.5 47.2 37.7 21.8 41.1 49.5 - MTL (SFT) 95.2 69.9 64.5 63.4 18.1 61.6 57.5 62.9 - SFT 76.1 55.8 46.8 58.5 20.2 62.2 58.2 54.0 -10.4 ReMax 87.6 71.4 51.6 62.4 16.8 33.3 54.1 53.9 -3.8 RLOO 94.0 73.7 48.9 62.7 42.1 40.5 55.3 59.6 -2.1 GRPO 93.0 74.8 51.8 65.9 38.4 41.3 54.2 60.0 -2.3
Table 2: General capabilities evaluation on MMMU, MMLU-Pro, and POPE benchmarks after training on downstream tasks. The best and second-best results are highlighted.
Benchmark Eval CoT Base SFT RFT SFT MTL (SFT) GRPO RLOO ReMax MMMU 52.1   40.1 ($\downarrow$12.0) 47.8 ($\downarrow$4.3) 54.2 ($\uparrow$2.1) 53.7 ($\uparrow$1.6) 48.7 ($\downarrow$3.4) 48.7 39.9 ($\downarrow$8.8) 48.1 ($\downarrow$0.6) 51.0 ($\uparrow$2.3) 46.8 ($\downarrow$1.9) 51.6 ($\uparrow$2.9) MMLU-Pro 47.5   30.6 ($\downarrow$16.9)   33.2 ($\downarrow$14.3) 48.3 ($\uparrow$0.8) 45.1 ($\downarrow$2.4)   35.4 ($\downarrow$12.1) 40.0 30.1 ($\downarrow$9.9) 32.9 ($\downarrow$7.1) 42.2 ($\uparrow$2.2) 39.7 ($\downarrow$0.3) 41.0 ($\uparrow$1.0) POPE 86.6 81.4 ($\downarrow$5.2) 84.9 ($\downarrow$1.7) 88.5 ($\uparrow$1.9) 88.2 ($\uparrow$1.6) 85.2 ($\downarrow$1.4) 82.2 81.5 ($\downarrow$0.7) 84.5 ($\uparrow$2.3) 84.0 ($\uparrow$1.8) 82.0 ($\downarrow$0.2) 87.2 ($\uparrow$5.0)
4.2 Finding 1: RFT Inherently Resists Catastrophic Forgetting
Our primary investigation focuses on the knowledge retention capabilities of SFT and RFT within a continual learning sequence. The results, summarized in Table 1, reveal a contrast between the two paradigms.
SFT suffers from catastrophic forgetting.
We observe that sequential SFT leads to a severe degradation of performance on previously learned tasks with a forgetting measure (FM) of -10.4%. For instance, performance on ScienceQA drops dramatically (95.2% $\rightarrow$ 76.1%) after completing the entire task sequence. The final average accuracy (AvgAcc) of 54.0% is substantially lower than the multi-task learning of SFT, which is the upper bound of 62.9%, confirming that SFT is highly susceptible to forgetting.
RFT preserves task knowledge and achieves MTL performance.
In contrast, all RFT methods demonstrate remarkable resilience against forgetting. As shown in Table 1, RFT methods exhibit very low forgetting measures, with GRPO achieving an FM of -2.3%. For example, GRPO maintains ScienceQA performance at 93.0% after learning all tasks, compared to its peak performance of 95.6%, which is a minimal drop compared to SFT. Among RFT methods, GRPO performs best, achieving a final AvgAcc of 60.0%, which is close to the upper bound of 62.9%. The model achieves this high performance without any explicit continual learning strategies, suggesting that the RFT paradigm is inherently robust for CPT.
4.3 Finding 2: RFT Protects and Enhances General Capabilities
Beyond task-specific knowledge, an ideal CPT process also requires preserving the model’s foundational, general-purpose abilities. We evaluated the models on general benchmarks to measure this effect. The results, presented in Table 2, highlight another critical advantage of RFT.
SFT harms general capabilities in both CL and MTL.
Our experiments reveal that SFT causes significant “base model degradation” (Liu et al., 2025a). SFT induces a severe performance drop of $\downarrow$16.9on the challenging MMLU-Pro benchmark (47.5% $\rightarrow$ 30.6%). Crucially, this is not merely an artifact of sequential learning; even multi-task SFT (MTL (SFT)), which trains on all data simultaneously, still causes a severe drop of $\downarrow$14.3on the same benchmark. A similar trend is evident on MMMU, where SFT and MTL (SFT) cause performance to decline by $\downarrow$12.0and $\downarrow$4.3respectively. This demonstrates that the SFT paradigm itself appears harmful to the model’s foundational capabilities.
RFT preserves and enhances general capabilities.
In contrast to the capability decay observed across all SFT methods, the RFT paradigm effectively safeguards the model’s general abilities. GRPO, in particular, often enhances these abilities. For instance, GRPO improves performance on MMMU by $\uparrow$2.1(52.1% $\rightarrow$ 54.2%). Crucially, RFT also improves model general capabilities, with GRPO improving the POPE score by $\uparrow$1.9(86.6% $\rightarrow$ 88.5%) and reducing the tendency for hallucination. This clear difference highlights that RFT is a more robust paradigm for continual post-training.
5 How Does RFT Mitigate Forgetting?
To investigate the mechanisms behind RFT’s remarkable stability, this section presents a series of ablation studies based on the popular and representative GRPO algorithm Shao et al. (2024).
Table 3: Downstream task performance for ablation models. We investigate the role of the KL term and CoT through variants of GRPO. indicates that the training process is unstable and requires multiple restarts from a previous checkpoint to achieve convergence.
Method SciQA TextVQA VizWiz GQA Geo. PathVQA sCLEVR AvgAcc SFT 76.1 55.8 46.8 58.5 20.2 62.2 58.2 54.0 GRPO 93.0 74.8 51.8 65.9 38.4 41.3 54.2 60.0 GRPO w/o KL 93.0 75.0 51.6 65.9  35.6  40.9  54.7 59.5 GRPO w/o CoT 94.7 74.7 63.8 65.9 23.8 38.2 54.4 59.4
Table 4: General capabilities evaluation for ablation models. Each benchmark is evaluated with and without CoT prompts to provide a comprehensive view.
Benchmark Eval CoT Base GRPO GRPO w/o CoT GRPO w/o KL MMMU 52.1 54.2 ($\uparrow$2.1) 51.8 ($\downarrow$0.3) 52.2 ($\uparrow$0.1) 48.7 51.0 ($\uparrow$2.3) 51.6 ($\uparrow$2.9) 49.2 ($\uparrow$0.5) MMLU-Pro 47.5 48.3 ($\uparrow$0.8) 48.9 ($\uparrow$1.4) 45.2 ($\downarrow$2.3) 40.0 42.2 ($\uparrow$2.2) 41.9 ($\uparrow$1.9) 42.3 ($\uparrow$2.3) POPE 86.6 88.5 ($\uparrow$1.9) 85.3 ($\downarrow$1.3)   74.2 ($\downarrow$12.4) 82.2 84.0 ($\uparrow$1.8) 88.7 ($\uparrow$6.5) 87.6 ($\uparrow$5.4)
5.1 The Roles of CoT and KL Penalty
We test two primary hypotheses: (1) The KL-divergence penalty, by regularizing policy updates, acts as a form of knowledge distillation (Li & Hoiem, 2017) that preserves past knowledge. (2) The complex reasoning structure of CoT builds more abstract and resilient knowledge representations, protecting them from being overwritten. Thus, we evaluate three GRPO variants against the SFT baseline: GRPO w/o KL: trained with CoT prompts but without the KL penalty term. GRPO w/o CoT: trained without CoT prompts, using direct question-answering format but retaining the KL penalty.
KL penalty is not the primary factor for preserving task-specific knowledge.
As shown in Table 3, removing the KL penalty (GRPO w/o KL) causes no degradation in performance on the continual learning sequence. The final average accuracy remains, demonstrating that the KL penalty is not the primary mechanism preventing task-specific catastrophic forgetting. However, it is crucial to note that the training process without the KL penalty exhibits significant instability in the later stages of the task sequence. These results are obtained after multiple attempts, re-initializing from the previous task’s checkpoint to achieve a convergent outcome, which suggests KL penalty plays a critical role in stabilizing the RFT process.
Figure 4: Loss landscapes illustration of SFT and RFT based continual post-training.
CoT is a performance booster, not a forgetting mitigator.
Our second hypothesis is also not supported by the data. The model trained without CoT (GRPO w/o CoT) still strongly resists forgetting, maintaining a high average accuracy across the task sequence (Table 3). In fact, it outperforms GRPO on VizWiz ( 63.8% vs. 51.8%). The general capabilities evaluation in Table 4 further confirms this conclusion. The GRPO w/o CoT model remains robust, and it achieves the highest score on the POPE benchmark (88.7%) when tested in non-CoT format evaluation. This demonstrates that while CoT can enhance performance on certain types of tasks, it is not the mechanism responsible for RFT’s resistance to catastrophic forgetting. Besides, as shown in Table 3, we observe that for GRPO w/o KL, using CoT during inference would lead to notable hallucination.
5.2 Theoretical Insight: Implicit Regularization from Reward Variance
To explore the difference between SFT and RFT, we analyze their gradient updates from the perspective of continual learning, revealing that RFT’s forgetting mitigation stems from an implicit regularization mechanism.
We formalize this by defining a measure of how much a gradient update interferes with knowledge from past tasks. Following foundational work in continual learning (Kirkpatrick et al., 2017), we use the Fisher Information Matrix (FIM) as a theoretical tool to quantify parameter sensitivity.
Definition 5.1 (Forgetting Risk).
Let $\mathcal{D}_{1:k-1}$ be the data from all previously learned tasks. The FIM is defined as $F_{k-1}\triangleq\mathbb{E}_{(x,a^{*})\sim\mathcal{D}_{1:k-1}}[\nabla_{\theta}\log\pi_{\theta}(a^{*}|x)(\nabla_{\theta}\log\pi_{\theta}(a^{*}|x))^{\top}]$. The forgetting risk of a gradient update $g$ for the current task $k$ is defined as its squared Mahalanobis norm with respect to the FIM of past tasks:
$\mathcal{R}(g)\triangleq g^{\top}F_{k-1}g.$ (9)
This risk measures the update’s magnitude in parameter subspaces critical for prior knowledge. Note that $F_{k-1}$ is a theoretical construct for our analysis and is not computed in practice.
For a single data point $(x_{k},a_{k}^{*})\in\mathcal{D}_{k}$, the SFT loss gradient is $g_{\text{SFT}}=-\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k})$. In contrast, the RFT policy gradient for a sampled response $a\sim\pi_{\theta}(\cdot|x_{k})$ is $g_{\text{RFT}}(a)=A(x_{k},a)\nabla_{\theta}\log\pi_{\theta}(a|x_{k})$, where $A(x_{k},a)$ is an advantage function ( $r(x_{k},a)-b(x_{k})$).
Theorem 5.2 (RFT’s Implicit Regularization Effect).
Consider a single update on task $k$ at parameters $\theta_{k-1}$. Let the rewards be normalized, $r(x_{k},a)\in[0,1]$. Under the technical assumptions specified in Appendix B, the expected forgetting risk of an RFT update is related to the SFT risk by:
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]=\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E},$ (10)
where $\mathcal{E}$ is an error term that is small when the policy is reasonably accurate. The term $\text{Var}[r(x_{k},a)]$ is bounded by $1/4$ for normalized rewards.
The full proof is provided in Appendix B. Theorem 5.2 reveals that the expected RFT risk is scaled by the variance of reward. When the model is uncertain about the correct answer for a given prompt, the reward variance is bounded, leading to an update in parameter subspaces important for old tasks, as shown in Figure 4.
5.3 RIF-RFT: Enhancing Stability and Efficiency of RFT
Our analysis in Section 5.2 reveals that RFT’s resilience to forgetting is based in a reward-variance-scaled regularization. However, this mechanism’s effectiveness relies on the model’s ability to generate responses that produce a meaningful reward signal. We identify a critical failure mode when the model is faced with what we term incompetent samples: training instances for which the current policy $\pi_{\theta}$ consistently fails to produce non-zero rewarded outputs. For such samples, the advantage estimates $A(x,a)$ are dominated by noise, leading to high-variance gradient estimates.
To address this challenge, we propose a simple yet effective method: Rollout-based Instance Filtering for RFT (RIF-RFT). The motivation, illustrated in Figure 4, is to prune the training data by identifying and discarding these incompetent samples before the RFT training. By filtering them out, RFT focuses its capacity on instances where it can receive a productive learning signal, stabilizing the regularization effect and improving efficiency.
The mechanism is formalized in Algorithm 1. For each instance in a new task’s dataset $\mathcal{D}_{k}$, we perform a small number of policy rollouts. If at least one of these rollouts produces a response with a reward greater than a minimal threshold $\tau$, we classify the instance and retain it in $\mathcal{D}_{k}^{\text{filt}}$. As shown in Table 5, RIF-RFT achieves competitive performance (57.5% of AvgAcc) and strong knowledge retention (-4.5% of FM), which is superior to SFT with a FM of -10.4%. Meanwhile, RIF-RFT use only 48.4% of training data. This demonstrates a compelling trade-off between efficiency and robustness.
Input: New task training set: $\mathcal{D}_{k}=\{(x_{i},y_{i}^{*})\}_{i=1}^{M}$; Current model policy: $\pi_{\theta}$; Number of rollouts per input: $N$; Reward threshold: $\tau$
Initialize filtered dataset: $\mathcal{D}_{k}^{\text{filt}}\leftarrow\emptyset$;
for each sample $(x_{i},y_{i}^{*})\in\mathcal{D}_{k}$ do
       Initialize $R_{\max}\leftarrow-\infty$;
       for $j=1$ to $N$ do
             Sample a response: $y_{ij}\sim\pi_{\theta}(\cdot\mid x_{i})$;
             Compute reward: $R(y_{ij})$;
             Update: $R_{\max}\leftarrow\max(R_{\max},R(y_{ij}))$;
      if $R_{\max}>\tau$ then
             Add $(x_{i},y_{i}^{*})$ to $\mathcal{D}_{k}^{\text{filt}}$;
Perform standard RFT on the filtered dataset $\mathcal{D}_{k}^{\text{filt}}$ to obtain $\pi_{\theta^{\prime}}$;
Output: Updated model $\pi_{\theta^{\prime}}$
Algorithm 1 Rollout-based Instance Filtering for RFT (RIF-RFT)
Table 5: Performance and data efficiency comparison of our proposed RIF-RFT.
Method SciQA TextVQA VizWiz GQA Geo. PathVQA sCLEVR AvgAcc FM SFT 76.1 55.8 46.8 58.5 20.2 62.2 58.2 54.0 -10.4 GRPO 93.0 74.8 51.8 65.9 38.4 41.3 54.2 60.0 -2.3 RIF-RFT 92.9 73.7 46.6 63.0 32.3 40.5 53.7 57.5 -4.5 Data Kept 81.4% 45.6% 42.1% 67.6% 37.2% 42.5% 52.3% - -
6 Conclusion
This work presents a comprehensive investigation into the role of the fundamental learning paradigm in continual post-training for MLLMs. Our central finding is that RFT naturally mitigates the catastrophic forgetting that plagues the standard SFT. Through extensive experiments, we demonstrate that while SFT leads to severe degradation of both previously learned task-specific skills and general capabilities, RFT paradigms inherently preserve those knowledge, achieving performance comparable to an offline multi-task learning setting. Our in-depth analysis reveals that this superiority is not due to explicit mechanisms like CoT or KL regularization, but is instead a result of the implicit regularization from reward-variance-weighted updates in RFT. Finally, we introduce RIF-RFT, an efficient instance filtering method that improves the stability and sample efficiency of RFT without compromising its robustness. This research suggests that RFT is not merely an alternative but a fundamentally more suitable paradigm for the continual and lifelong adaptation of foundation models.
Appendix A Dataset Information
NoCoT Prompt Template
Question:
<image>What is the probability that a Nile tilapia fish...
A. 2/4   B. 3/4 ... E. 4/4
You MUST provide the final answer directly.
Answer:
E
CoT Prompt Template
Question:
<image>What is the probability that a Nile tilapia fish...
A. 2/4   B. 3/4 ... E. 4/4
You FIRST think about the reasoning process...
The reasoning process MUST BE enclosed within
<think> </think> tags.
[…]
Answer:
<think>
The Punnett square shows...
Therefore, the probability is 4/4. The correct answer is E.
</think>
[…]
Figure 5: Example prompt templates w/o and w/ CoT.
Multimodal Datasets for Continual Post-Training and Evaluation.
Our study utilized a diverse suite of vision-language datasets for both model training and comprehensive evaluation of various multimodal capabilities, along with specialized benchmarks to assess knowledge retention and nuanced multimodal challenges. Here is a brief introduction to these datasets:
Multimodal Datasets for Continual Post-Training:
  • ScienceQA (Saikh et al., 2022) presents multimodal science questions requiring complex reasoning over diagrams, text, and general knowledge.
  • TextVQA (Singh et al., 2019) focuses on questions that necessitate reading and inferring from text embedded within images.
  • VizWiz (Gurari et al., 2018) comprises real-world image-based questions posed by visually impaired individuals, often involving ambiguity.
  • GQA (Hudson & Manning, 2019) is designed for compositional question answering over real-world images with a strong emphasis on spatial understanding and object relationships.
  • Geometry3K (Lu et al., 2021): This subset of MathVista (Lu et al., 2024) comprises multi-choice geometry problems equipped with dense annotations in formal language for both diagrams and text, specifically designed to evaluate complex geometric reasoning skills.
  • PathVQA (He et al., 2020) provides medical visual question answering on pathology images that demand specialized domain knowledge.
  • Super-CLEVR (Li et al., 2023b) is a synthetic dataset crafted to rigorously test complex relational and logical reasoning.
Benchmarks for General Knowledge Evaluation:
  • MMMU (Yue et al., 2024) is comprehensive benchmark comprising 11.5K college-level, multi-discipline multimodal tasks with diverse image types, demanding deliberate reasoning.
  • MMLU-Pro (Wang et al., 2024b) is an enhanced benchmark designed for more discriminative evaluation of large language models, featuring more challenging and reasoning-focused questions with ten multiple-choice options, sourced from various academic and STEM fields.
  • POPE (Li et al., 2023a) is a benchmark introduced to systematically investigate and assess object hallucination in vision-language large models through an improved polling-based query method.
Appendix B Proof and Technical Details for Theorem 5.2
We provide the detailed derivation for Theorem 5.2, which establishes the relationship between the forgetting risks of RFT and SFT.
Theorem 5.2.
Let the rewards be normalized, $r(x_{k},a)\in[0,1]$. Under Assumption B.2, the expected forgetting risk of an RFT update is related to the SFT risk by:
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]=\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E},$ (11)
where $\mathcal{E}$ is an error term characterized in the proof.
Definition B.1 (Importance-Weighted Score Norm (IWSN)).
For a response $a$, we define its IWSN as the squared norm of its score function, weighted by the FIM of past tasks:
$I(a)\triangleq(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k})).$ (12)
Assumption B.2 (Technical Assumptions).
Our analysis relies on the following two technical assumptions for a given data point $x_{k}$ and parameters $\theta_{k-1}$:
  1. 1.
    Bounded Covariance. The covariance between the squared advantage and the IWSN is bounded: $\text{Cov}(A(a)^{2},I(a))=\epsilon_{1}$, where $\epsilon_{1}$ is a small error term. This implies that the magnitude of an advantage signal is not strongly correlated with the gradient’s impact on prior tasks.
  2. 2.
    Centered Policy Expectation. The expected IWSN under the current policy is close to the IWSN of the ground-truth response: $\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[I(a)]-I(a_{k}^{*})=\delta$, where $\delta$ is another small error term. This holds when the policy $\pi_{\theta_{k-1}}$ generates responses that, on average, have a similar gradient geometry to the ground-truth response with respect to past tasks.
Lemma B.3 (Variance of Advantage).
For policy gradient methods, using the reward baseline $b(x_{k})=\mathbb{E}_{a\sim\pi_{\theta}}[r(x_{k},a)]$ minimizes the variance of the gradient estimator. With this optimal baseline, the expected squared advantage equals the reward variance:
$\mathbb{E}_{a\sim\pi_{\theta}}[A(x_{k},a)^{2}]=\text{Var}_{a\sim\pi_{\theta}}[r(x_{k},a)].$ (13)
Proof.
By definition, $A(x_{k},a)=r(x_{k},a)-b(x_{k})$. We have:
$\mathbb{E}[A^{2}]$ $=\mathbb{E}[(r-b)^{2}]=\mathbb{E}[r^{2}]-2b\mathbb{E}[r]+b^{2}.$ (14)
Since $b=\mathbb{E}[r]$, this simplifies to $\mathbb{E}[r^{2}]-2(\mathbb{E}[r])^{2}+(\mathbb{E}[r])^{2}=\mathbb{E}[r^{2}]-(\mathbb{E}[r])^{2}=\text{Var}[r]$. ∎
Proof of Theorem 5.2.
Let us analyze the forgetting risks at parameters $\theta=\theta_{k-1}$ for a single data point $(x_{k},a_{k}^{*})$. The analysis proceeds in three steps.
Step 1: SFT Forgetting Risk.
The SFT loss gradient is $g_{\text{SFT}}=-\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k})$. Its forgetting risk is deterministic:
$\mathcal{R}(g_{\text{SFT}})=g_{\text{SFT}}^{\top}F_{k-1}g_{\text{SFT}}=(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))=I(a_{k}^{*}).$ (15)
Step 2: Expected RFT Forgetting Risk.
The RFT gradient for a sampled response $a\sim\pi_{\theta}(\cdot|x_{k})$ is $g_{\text{RFT}}(a)=A(x_{k},a)\nabla_{\theta}\log\pi_{\theta}(a|x_{k})$. We compute the expectation of its forgetting risk:
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}_{a\sim\pi_{\theta}}\left[(g_{\text{RFT}}(a))^{\top}F_{k-1}(g_{\text{RFT}}(a))\right]$ (16)
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))\right]$ (17)
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot I(a)\right].$ (18)
Step 3: Relating RFT and SFT Risks.
We use the covariance identity $\mathbb{E}[XY]=\mathbb{E}[X]\mathbb{E}[Y]+\text{Cov}(X,Y)$ to decompose Eq. 18:
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}[A(a)^{2}]\cdot\mathbb{E}[I(a)]+\text{Cov}(A(a)^{2},I(a))$ (19)
[…]
$=\text{Var}[r(x_{k},a)]\cdot I(a_{k}^{*})+\underbrace{\text{Var}[r(x_{k},a)]\delta+\epsilon_{1}}_{\text{Error Term }\mathcal{E}}$ (22)
$=\text{Var}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E}.$ (23)
This completes the proof. The error term $\mathcal{E}$ is small when both $\epsilon_{1}$ and $\delta$ are small, which holds when the policy is reasonably well-behaved and the advantage signal is not strongly correlated with parameter importance. The primary relationship demonstrates that the expected forgetting risk of RFT is modulated by the reward variance, providing an implicit regularization against catastrophic forgetting. ∎